Suppose that Gyp Sum Industries currently has the balance sheet shown below, and that sales for the year just ended were $10.3 million. The firm also has a profit margin of 25 percent, a retention ratio of 30 percent, and expects sales of $8.3 million next year. Assets Liabilities and Equity $2,189,000 Current liabilities Current assets Fixed assets $1,816,920 1,650,000 3,022,080 $6,489, 000 4,300,000 Long-term debt Equity Total assets $6,489,000 Total liabilities and equity If all assets and current liabilities are expected to shrink with sales, what amount of additional funds will Gyp Sum need from external sources to fund the expected growth? (Enter your answer in dollars not in millions. Negative amount should be indicated by a minus sign.) Additional funds needed
Answer #1

Return on Equity = Net Income/ Equity = ( 10.3 × 25%)/3022080= 85.21%

Growth Rate = Return on Equity × Retention Ratio

Growth Rate = 85.21% × (30/100) = 25.562 %

Expected Profit for the next year = 8.3 × 0.25 = 2.075 million

ROE should be 85.21%. Equity will be 2.075/.8521

2435160.20

Which is 586920 less than 3022080.

-586920 Answer